OpenStudy (anonymous):

solve by elimination. please show work, and state point of intersection. 6x-y=21 and 4x+3y=17

OpenStudy (anonymous):

3(6x-y=21) 4x+3y=17 18x - 3y = 63 4x + 3y = 17 22x = 80 x = 40/11 Substitute back for y. (40/11, 9/11)

OpenStudy (anonymous):

that doesnt work.

OpenStudy (anonymous):

The universe has SPOKEN :D.. Its proper!

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

I rechecked my work, it is correct.

OpenStudy (anonymous):

how if i plug it back in the equation it doesnt work!